OpenStudy (anonymous):

32 is 40% of what number? I didnt have math for 2 years now and forgot everything? I need help with this problem. Thank YOu. With explanation please!!:)

OpenStudy (reemii):

40% of any number \(X\) can be simply computed as \(\frac{40}{100}x\). Now you ask that \(\frac{40}{100}\text{some number \(x\)}=32\). You just have to solve \(\frac{40}{100}x=32\).

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (reemii):

'welcome :)